### v1

# Careful finite element simulations of cold rolling with accurate through-thickness resolution and prediction of residual stress

Abstract
In this paper, we carefully develop a finite element (FE) model that gives accurate through-thickness predictions of stress and strain distributions during the steady-state cold rolling process. These through-thickness predictions unveil an oscillatory pattern that is shown to have important consequences for residual stress in the rolled sheet. We believe this is the first time that through-thickness FE results have been accurately validated by comparison to non-FE results, in this case by comparison to a recent analytical model of through-thickness variation in cold rolling. While we use here the ABAQUS commercially available FE software, our observations are relevant to all FE simulations of cold rolling. Care is taken by considering both convergence in number of elements through thickness, convergence to a steady state, and the avoidance of other numerical artefacts such as shear locking and hourglassing. We find that previous FE models of cold rolling are usually woefully under-resolved through-thickness; e.g. using 10 elements through-thickness, while we require 60 here for convergence. Convergence of roll force and roll torque, used in previous studies to validate models, are shown to be poor indicators of through-thickness convergence. We also show that the through-thickness oscillatory pattern may have important consequences for predicting curvature during asymmetric rolling.
keywords:
finite element , plasticity , elasticity , cold rolling , residual stress.
journal: Int. J. Mech. Sci.
[…]
[macsi]organization=MACSI, Department of Mathematics & Statistics, addressline=University of Limerick, city=Limerick, postcode=V94 T9PX, country=Ireland
[…]
[csis]organization=Computer Science and Information Systems (CSIS), addressline=University of Limerick, city=Limerick, postcode=V94 T9PX, country=Ireland
[…]
[lero]organization=Lero, The Science Foundation Ireland Centre for Software Research, addressline=University of Limerick, city=Limerick, postcode=V94 T9PX, country=Ireland
[…]
[wmi]organization=Mathematics Institute, addressline=University of Warwick, city=Coventry, postcode=CV4 7AL, country=UK
[…]
[cued]organization=Department of Engineering, addressline=University of Cambridge, city=Cambridge, postcode=CB2 1PZ, country=UK
[…]
[ted]organization=Mechanical Engineering Department, addressline=TED University, city=Ankara, country=Turkey
[…]
[wmg]organization=WMG, addressline=University of Warwick, city=Coventry, postcode=CV4 7AL, country=UK
1 Introduction
Energy used in industrial activity produces 40% of global $\mathrm{CO}_{2}$ emissions, the majority of which are caused by the manufacture of primary materials such as steel and aluminium [1]. Steel manufacturing is particularly problematic when one considers that crude steel production grew by approximately 900% between 1950 and 2023 [2], and that global steel demand is expected to grow by another 2.9% by 2025 [3]. The underlying mechanics of metal forming processes are not yet fully understood, and hence the problem of efficiently modifying the tool path based on the measured error is still unsolved [4]. As a result of these misunderstandings, liquid metal must be overproduced to ensure final product quality. For manufacturers this increases costs and reduces productivity [5] but, as liquid metal production is the most energy and $\mathrm{CO}_{2}$-intensive processing step, these knowledge gaps also have significant implications for sustainability. Reducing liquid metal demand through process improvements could have a significant impact on energy use and $\mathrm{CO}_{2}$ emissions [6]. In this work we will focus on cold rolling, a common steel manufacturing method.
In cold rolling, a sheet of metal, whose temperature is below the material recrystallisation temperature, is gradually fed between two rotating rollers. The rollers are held at a fixed separation that is less than the initial thickness of the sheet and act to permanently deform the sheet so that its thickness is reduced as illustrated in Figure 1.
Figure 1: Schematic of the metal rolling process. The sheet moves from left to right and its thickness is reduced by the roller. Symmetry about the centre of the sheet is assumed.
The most common method of assessing cold rolling is through the application of mathematical models, which can be used to optimise system design and tighten process specification [1, 7]. Most traditional slab-type mathematical models are based on the classical Orowan [8] approach. In general, slab models assume that the stresses are vertically homogeneous through the thickness of the sheet, that each vertical slab of material remains vertical and that the sheet does not stick to the rollers throughout the rolling process. While this very simple mathematical description can yield reasonably accurate predictions of the roll force and torque required for a desired thickness reduction, they disregard through-thickness effects [9, 10]. These through-thickness distributions are significant for downstream properties such as residual stress which in turn informs the overall quality of the final product. It is therefore desirable to enhance the understanding of the through-thickness stress and strain in the sheet.
Through-thickness variations in stress and strain can be attained through more advanced modelling techniques such as finite element (FE) analysis. The FE method can give high-resolution predictions of stresses and strains in the roll gap. In contrast to existing analytical models, FE models also predict stress distributions outside the roll gap. Recent FE simulations of cold rolling by the current authors [11] show oscillatory stress and strain patterns. These distributions appear to have important consequences for cold-rolling residual stresses [12, 13, 11], while slab models fail to capture these variations. Of course, the FE method is still an approximate solution technique with its own limitations, such as slow computational time. This makes the FE method unsuitable for real-time process control [14]. Convergence issues for extreme parameter values also limit the range of possible explorations that can be considered during FE simulations. For example, elasticity is often neglected by assuming an infinite Young’s modulus, $E$, in analytical modelling for simplification, but this is impossible to reproduce in certain elastic-plastic FE softwares such as ABAQUS. Analytical approaches such as asymptotics-based models have the potential to be used in online control, assuming they can output sufficiently accurate results. Although not directly useful for online control, FE outputs can be used as a benchmark for these quicker-to-compute analytical models to ensure that the analytical modelling gives satisfactory outputs. The FE method allows for a range of parameter regimes to be explored at a relatively low cost compared to experimental results [10]. To act as a benchmark for through-thickness predictions, the FE model used needs a sufficiently fine mesh, among other model settings, which we will see later in this work. The overarching aim of this paper is to provide a reliable FE model that accurately captures the through-thickness stress and strain variations arising in the sheet during rolling, which can guide the development of simpler, faster mathematical models. As we will see, such a reliable FE model is absent from existing literature.
In Section 2, we provide a brief overview of existing analytical and FE cold-rolling models that consider through-thickness stress and strain variations, while an in-depth description of the ABAQUS/Standard and ABAQUS/Explicit FE formulations may be found in the supplementary material. In Section 3 we provide details of the implicit and explicit FE cold rolling models. In Section 4 we conduct a mesh study and compare implicit and explicit in terms of their ability to accurately predict through-thickness stress variations. We show that roll force and torque are poor determinants of the optimal mesh size and of the steady state condition. In Section 5, we compare the simulated FE results with a new analytical model from Erfanian et al. [13]. We also show that the residual stress in the sheet is related to the mechanics inside the roll gap. In Section 6 we draw conclusions and briefly discuss the future plans for our FE simulations.
2 Literature review
Here we provide an overview of some existing analytical and FE rolling models that consider through-thickness variations in their analyses, and demonstrate their limitations.
A range of commerical FE packages are available for modelling metal deformation, such as Ansys, DEFORM, LS-Dyna, MSC Marc and ABAQUS. FE simulations reported in this work were conducted using ABAQUS 2021 [15]. Within the ABAQUS package there are two numerical solution types: implicit and explicit. For the cold rolling process it is not immediately clear which solution type is optimal as each type offers different advantages and disadvantages. The explicit solver is often the preferred choice for modelling metal forming processes due to its computational speed and robust handling of complex contact conditions [16, 17, 18, 19]. A key disadvantage of choosing the explicit solver is the absence of equilibrium convergence enforcement which can lead to error accumulation [20]. The potential for explicit solvers to accumulate error implies that implicit solvers provide stronger results [21]. Implicit solvers however typically consume more computational resources and can struggle to handle complex contact [19]. In Section 3 we will assess both implicit and explicit solvers for modelling cold rolling, with the specific aim of accurately capturing through-thickness variations of stress and strain quantities.
Asymptotic analysis is useful for modelling physical processes where the smallness (or largeness) of a fundamental process parameter can be exploited [22]. This type of analysis is particularly useful for cold rolling, where the roll-gap aspect ratio is typically large [14], i.e., $\varepsilon=h_{0}/L\ll 1$, where $h_{0}$ is the initial half-thickness of the sheet and $L\approx\sqrt{2R\Delta h}$ is the horizontal length of the roll gap (see Figure 1). Domanti and McElwain [23] assume a small reduction in the sheet’s thickness (i.e., $r=\Delta h/h_{0}\ll 1$) along with the large-aspect-ratio assumption. These assumptions enable an asymptotic analysis to be carried out to obtain accurate solutions in the limits as $\varepsilon\rightarrow 0$ and $r\rightarrow 0$. However, these assumptions limit the range of validity of the analytical model. Erfanian et al. [13] relax the assumption that the reduction of the sheet thickness is small and also resolve behaviour over two distinct length scales associated with $h_{0}$ and $L$. By contrast, Domanti and McElwain [23] and other authors [see, for example, 14] consider variations on a length scale of $L$ only. The multiple-scales analysis of Erfanian et al. [13] recovers slab theory at leading order, and at higher orders predicts through-thickness oscillations in stress and velocity in the roll gap. These through-thickness variations have also been captured by the current authors in recent FE simulations [11]. The Erfanian et al. [13] model is limited to rigid-perfectly-plastic materials, thus elastic and strain-hardening effects are neglected. FE models can integrate elastic-plastic material behaviour which captures elasticity and strain hardening. Comparing FE outputs to analytical models such as the Erfanian et al. [13] model can give an understanding of what mathematical assumptions are valid or invalid.
Table 1: Summary of FE models in the literature that consider through-thickness variations. Note that $2N_{e}$ represents the number of elements through the full thickness of the sheet, PEEQ is plastic equivalent strain and the material models are elastic-plastic (EP), rigid-plastic (RP), elasto viscoplastic (EVP) and elastoplastic (EoP). The * indicates that oscillations are visible in certain FE results, but are not directly discussed in the paper.
Reference 2D or 3D $2N_{e}$ Element type Material model Through-thickness variations Symmetric system Oscillatory pattern
Current work 2D 10–80 Plane-strain 4-node, reduced-integration EP Shear, von Mises, PEEQ, PEEQ rate
[24] 2D 10 RP Mesh distortion, relative velocities, normal/shear strain rates, PEEQ
[25] 2D 12 Six-node triangular EVP Effective stress
[26] 2D 10 Isoparametric quad full integration RP Mesh distortion, PEEQ rate, normal stress
[27] 3D 8–10 8-node brick EP Velocity flow patterns
[28] 2D 8 Quad, reduced-integration EP Stress components
[29] 3D 6 EP Mesh distortion
[30] 2D 14 Quad RP PEEQ rate, pressure
[31] 3D 3 8-node brick EP Incremental strain components
[32] 2D 20 Quad EP Mesh distortion, PEEQ
[33] 2D 8 RP PEEQ, PEEQ rate Asymmetric
[34] 2D 18 4-node EoP Stress (centre & surface only)
[35] 2D 18 Quad EVP Mesh distortion, PEEQ, horizontal stress Asymmetric
[36] 3D 10 8-node brick, reduced-integration EP Effective stress, PEEQ
[37] 2D 8 Rectangular RP PEEQ
[38] 2D 16–18 Plane-strain 4-node VP Temperature, strain, strain rate, effective stress
[9] 2D Isoparametric quad elements EP PEEQ
[39] 3D 10 8-node brick reduced-integration EP Vertical, shear stress components
[40] 3D 12 8-node brick reduced-integration temperature-coupled EP Normal stress, pressure, displacements (edge only)
[41] 2D 16 Square isoparametric RP PEEQ, displacements, velocities
[42] 2D 36 3-noded triangular RP PEEQ, temperature
[43] Both 2D: 18-30, 3D: 11 2D: Plane-strain 4-node reduced-integration, 3D: 8-node reduced-integration EP Stress, velocity components Symmetric and asymmetric
[44] 2D 20 4-node isoparametric EP Temperature, heat flux
[45] 2D 16 4-node square elements EP Longitudinal, shear and effective strain, longitudinal and von Mises stress
[11] 2D 12–100 Plane-strain 4-node reduced-integration EP Vertical velocity, von Mises
Table 1 shows the common through-thickness variations from existing FE models. As shown in Table 1, comparing through-thickness variations from various works is difficult given the distinct differences in modelling criteria and assumptions (e.g., 2D or 3D simulation, number of elements through-thickness ($2N_{e}$), element type, etc.) made by authors. The implication of this lack of comparability means that there is no clear mechanism for novices to recognise what constitutes a “good” FE rolling simulation. While several authors report shear stress along the roller-sheet interface [16, 46, 14], Table 1 shows that few [28, 39, 43] provide the through-thickness shear stress variation. Given the potential implications of oscillating shear stress on downstream quantities such as residual stress [12, 11], a “good” FE model should be capable of accurately predicting the oscillatory shear stress (and other quantities) through the thickness of the sheet.
[…]
Olaogun et al. [44] show a similar oscillatory pattern in their heat flux contour plots. The maximum values of heat flux lie along a path that is similar to the path of the lines of maximum shear. Heat energy is dissipated during plastic flow [47], and plastic flow takes place in metals predominantly by shearing [48]. We therefore expect that a close relationship exists between the heat flux oscillations observed by Olaogun et al. [44] and the shear stress oscillations that exist during rolling. However, Olaogun et al. [44] do not present shear stress results since heat transfer during cold rolling is the main topic of this paper.
Finally, many other authors have illustrated through-thickness variation but only via quantities that are not significantly oscillatory, such as longitudinal stress, effective stress or equivalent strain [25, 26, 32, 33, 35, 36, 37, 9]. Others report through-thickness variations via mesh distortion [26, 29, 32]. Tadić et al. [45] study the longitudinal residual stress induced by cold rolling through-thickness inhomogeneities via FE analysis, and state that it is “imperative” to determine the through-thickness stress distributions during cold rolling to accurately predict and control the impact of longitudinal residual stress on the appearance and properties of cold-rolled strips. However, Tadić et al. [45] select their mesh density ($N_{e}=8$) by satisfying convergence conditions concerning the mean value of the contact stress. We will see in subsection 4.3 that accurate measurements of such surface quantities does not imply accurate through-thickness predictions, and that $N_{e}=8$ is far too coarse, especially when through-thickness predictions are desired. Li and Kobayashi [24] present contour plots of shear strain rate at a low resolution inhibiting any oscillatory pattern observation. Liu et al. [31] presented cross-sectional profiles of strain components from three-dimensional simulations which contain hints of oscillations but these patterns are not commented on. The results of Misovic et al. [41] for displacement and velocity also hint at a possibility of oscillations, but are only presented at the sheet centre and surface.
It is clear that a reliable FE model, one capable of accurately predicting through-thickness stress and strain variations, is absent from existing literature. We aim to provide such a model, which can be used to guide and compare with faster analytical models.
3 Simulation details
In this section we give a complete description of the FE models used in this work. We include key information about the geometry, material, dynamics, mesh, contact and computer precision settings.
3.1 Model geometry
Since the metal sheet’s width is much larger than its thickness in cold rolling in general, we consider an idealised two-dimensional geometry by assuming plane-strain conditions apply [46]. Readers are referred to Figure 1 for a diagram of the model geometry. We consider a rigid roller of radius $R=257.45\,$\mathrm{m}\mathrm{m}$$, rolling a sheet of initial half-thickness $h_{0}=2\,$\mathrm{m}\mathrm{m}$$, and impose an approximate half-thickness reduction of $\Delta h=0.5\,$\mathrm{m}\mathrm{m}$$. This corresponds to a roll-gap aspect ratio of $\varepsilon=h_{0}/L=0.125$, where $L\approx 16.05\,$\mathrm{m}\mathrm{m}$$ is the approximate horizontal length of the roll gap.
3.2 Material properties
We consider a sheet of elastic-plastic mild steel (grade DC04), a common material used in cold metal forming [49]. The DC04 material deforms elastically up to an initial yield stress of $Y=477.2\,$\mathrm{M}\mathrm{P}\mathrm{a}$$, and the yield stress increases to $650.25\,$\mathrm{M}\mathrm{P}\mathrm{a}$$, at a true plastic strain of 1.1, as a result of strain hardening. The Young’s modulus and Poisson’s ratio are taken to be $E=206.3\,$\mathrm{G}\mathrm{P}\mathrm{a}$$ and 0.3 respectively. For all explicit analyses, the density of the material is defined as $\rho=7831.3\,$\mathrm{k}\mathrm{g}\,\mathrm{m}^{-3}$$.
3.3 General FE model conditions
The model consists of two steps: the bite step where contact between the roller and the sheet is initialised, and the rolling step, where the sheet is horizontally displaced due to rotation of the roller. To ensure contact convergence between the sheet and the roller the bite step occurs over a time increment of $1\,$\mathrm{s}$$. The rolling step occurs over a time increment of $0.1\,$\mathrm{s}$$. This is enough time for approximately 8 roll-gap lengths to be rolled when $R=257.45\,$\mathrm{m}\mathrm{m}$$. The step type is dependent on the solver implemented (i.e. implicit or explicit), and readers are referred to Table 3 for the specific step types used in simulations discussed here.
The sheet’s horizontal centre line (i.e., the symmetry axis in Figure 1) is constrained vertically to enforce the expected symmetry. This modelling choice reduces the computational intensity of the problem, and in turn facilitates the use of a finer mesh. The centre of the roller is defined as a reference point and boundary conditions, depicting the movement of the roller, are applied to this location. Initially, the roller and the sheet are not in contact, but are set up such that the roller is positioned above the sheet. In the bite step the roller is slowly translated vertically to indent the sheet, but no rotation of the roller occurs in the bite step. A bite step is used here, instead of starting the strip in front of the roll gap and feeding it in in some way, as the bite-then-roll simulations showed fewer initial contact issues and converged more quickly to steady state (as also described by Minton [43]). The bite step is performed via a velocity boundary condition at the roller’s reference point. In the rolling step, the roller reference point is fixed so its only degree of freedom is rotation in the rolling direction, and the roller is given a tangential velocity of $R\Omega=1287.25\,$\mathrm{m}\mathrm{m}\,\mathrm{s}^{-1}$$. Note that the above boundary conditions are used for all analyses below unless otherwise stated.
As the formulation differs for implicit and explicit simulations the model specifications understandably diverge. Table 3 provides model settings unique to implicit and explicit simulations.
Table 3: Unique FE model settings for the ABAQUS/Standard and ABAQUS/Explicit rolling models used in this work.
Feature ABAQUS/Standard ABAQUS/Explicit
Density N/a $7831.3\,$\mathrm{k}\mathrm{g}\,\mathrm{m}^{-3}$$
Step type Static, general Dynamic, explicit
Amplitude curve Instantaneous Smooth step
Contact tracking State-based N/a
Contact discretisation Surface-to-surface N/a
Mechanical constraint formulation N/a Kinematic contact
Element types considered CPE4/CPE4I/CPE4R CPE4R
ABAQUS/Explicit precision N/a Double, analysis & packager
The amplitude curve information given in Table 3 relates to the application of loading on the roller during the rolling step. Implicit simulations detailed here apply an instantaneous $\Omega=5~{}$\mathrm{r}\mathrm{a}\mathrm{d}\,\mathrm{s}^{-1}$$ while explicit simulations smoothly ramp loading to $\Omega=5~{}$\mathrm{r}\mathrm{a}\mathrm{d}\,\mathrm{s}^{-1}$$ over a fraction of the rolling step.
In our analyses we compare results from the four-node, plane-strain elements. As detailed in Table 3, implicit analyses include comparisons of full integration (CPE4), reduced integration (CPE4R), and incompatible (CPE4I) type elements. Since full integration (CPE4) and incompatible type (CPE4I) elements are not offered by ABAQUS/Explicit only reduced integration (CPE4R) elements are considered. Meshing is applied using a global seed formulation with a structured meshing technique. Varying levels of mesh density are considered and the optimal mesh is discussed in detail in Section 4.
3.4 Contact conditions
The roller, as an analytically rigid geometry, is defined as a master surface for contact between the roller and the sheet. The surface of the sheet in contact with the roller is defined as the slave (i.e. deformable) surface. Hard pressure-overclosure contact is defined between the roller and the sheet [15]. For cold-rolling processes we expect significant relative motion between the rigid roller and the deformable sheet. To account for relative motion between the roller and sheet surfaces, a finite sliding formulation is adopted. Friction between the roller and the sheet is defined through the isotropic Coulomb friction model. Coulomb friction is regularly used in FE simulations of rolling (e.g., [24, 26, 33, 37, 43]). For industrial cold rolling mills the friction coefficient typically lies in the range of $\mu=0.02-0.15$ [50]. FE models in this work enforce a friction coefficient of $\mu=0.1$.
3.4.1 Implicit analyses
Implicit simulations employ multiple types of contact-tracking algorithms. For finite-sliding contact interactions between analytically rigid and deformable surfaces the state-based algorithm is recommended [15]. The state-based contact algorithm evaluates contact interactions based on the current state of the simulation, together with geometric information. Contact is discretised using the surface-to-surface method.
3.4.2 Explicit analyses
Mechanical constraints in explicit analyses are described using the default kinematic contact method.
3.5 Transient effects and job precision
Our focus in this paper is on steady-state behaviour, so we aim to minimize transient effects during the dynamic explicit analysis.
3.5.1 Implicit analyses
To ensure computational precision full nodal output was requested for all simulations in this work. This is equivalent to double precision in explicit simulations.
3.5.2 Explicit analyses
To minimise transient effects loading in explicit simulations was applied using the smooth amplitude method. Similarly mass scaling and/or increased loading rates were not applied. To further increase computational precision explicit simulations were conducted using double precision for both the packager and analyses. This is equivalent to the full nodal precision requested in implicit analyses.
4 FE model refinement
Here we investigate the effect of element type, solver types (i.e., implicit and explicit) and mesh size on through-thickness stress and strain. Through careful consideration of these variables we obtain a reliable FE model for the rolling process. We then describe the process of ensuring that the simulated rolling process is in a steady state.
4.1 Element type
In this work, first-order plane-strain elements are employed to discretise the sheet. These elements provide sufficient accuracy and are less computationally-expensive compared to second-order elements [51]. Others [16, 34, 52, 38, 44] similarly used first-order elements to model rolling. As previously discussed (see Section 3.3), ABAQUS limits the element type depending on solver implementation. The element types applied in this work are outlined in Table 3 which differentiates by solver implementation. In the following sections we discuss the benefits (and potential drawbacks) of each element type to demonstrate the effect of element selection on FE model output.
4.1.1 Implicit analyses
The material response is evaluated by Gaussian quadrature at each element integration point [15]. Plane strain four-node (CPE4) elements contain a total of four integration points. The number of nodes and/or number of integration points correlates to the degree of freedom any given element can exhibit. Reduced integration elements, such as CPE4R elements, use just a single integration point in the centre of the element (compared to the four integration points in a fully integrated element). Therefore, in principle, CPE4 elements are more accurate than the CPE4R elements. However, in the presence of bending the inherent linear geometry of CPE4 elements can give rise to large, artificial shear strains. This problematic behaviour is often termed “shear locking”  [53, 54]. Shear locking is a purely computational behaviour and is not representative of physical processes. FE models exhibiting shear locking provide erroneous displacement and stress values [53].
Figure 2: Shear stress evaluated at the integration points (i.e., no nodal averaging) for a rolling simulation using CPE4 elements. There are significant jumps in shear stress across various elements, creating a checkerboard pattern, which is typical of shear locking behaviour.
Figure 2 shows shear stress evaluated at the integration points of CPE4 elements during a rolling simulation conducted as outlined in the previous sections. In Figure 2 a clear checkerboard pattern is observed, with more extreme shear stress values observed at the corners of elements. This is typical of shear locking behaviour [15] and implies that CPE4 elements are not suitable for this test case, hence this element type is not considered for the remainder of this work.
Unlike CPE4 elements, reduced integration (CPE4R) and “incompatible” (CPE4I) elements do not suffer from shear-locking behaviour. Low-order elements lack the necessary shape functions to describe bending. Incompatible elements (CPE4I) artificially add the required shape functions to adequately describe this behaviour [55, 56]. Simulations conducted using CPE4R and CPE4I elements do not exhibit the checkerboard pattern observed in Figure 2. This indicates that both reduced-integration and incompatible elements are sufficient for removing shear locking behaviour. However, incompatible elements are slow to convergence in simulations with large compressive strains [15]. This severely restricts the mesh size that can be employed with the CPE4I element type (see Section 3 of the supplementary information provided, where the performance of CPE4I and CPE4R elements are compared), and hence CPE4I elements are not considered for the remainder of this work.
4.1.2 Implicit and explicit analyses
[…]
4.2 Limitations of ABAQUS/Explicit
We now consider results from ABAQUS/Explicit simulations. Mass scaling (or increased loading rate) was not used in any explicit simulation in this paper. Given that the kinetic-to-internal energy ratio remains in the desired range (5-10%), we consider inertia effects to be insignificant.
Figure 3: Contour plots showing von Mises stress ($\mathrm{M}\mathrm{P}\mathrm{a}$) from explicit simulations with (a) $N_{e}=5$ CPE4R elements and (b) $N_{e}=40$ CPE4R elements. Both simulations have roller radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
As mesh density increases, it is expected that the calculation accuracy should, in general, improve [61]. However, Figure 3 shows that increasing the mesh density increases noise. Figure 3 compares the von Mises stress for two explicit simulations with mesh densities of $N_{e}=5$ (Figure 3a) and $N_{e}=40$ (Figure 3b) elements through the half-thickness, respectively. The most notable difference between the plots is the smoothness of the solution. This result indicates that the solution diverges as mesh density increases. In particular, we note the noisy oscillatory behaviour near the surface of the sheet towards the entrance to the roll gap, $x/L=0$. Flanagan et al. [11] successfully removed noise by implementing additional stiffness-proportional damping. The damping approach however increases the required computational time [700 CPU hours compared to 10 CPU hours without damping for a particular rolling simulation conducted by 11]. Given the divergent nature of the explicit results and considering the lack of equilibrium checks we propose that implicit simulations outperform explicit simulations for this case study. Therefore the remainder of this work will focus on simulation results from implicit simulations only.
4.3 Mesh sensitivity study
In subsection 4.1 we demonstrated that CPE4R elements are the preferred element type for this simulation case study. We also demonstrated (subsection 4.2) that explicit analyses are subject to increasing noise with increasing mesh density. In this section we perform a convergence study for implicit simulations that results in the optimum trade-off between accuracy and computational intensity.
Figure 4: Time-history plots of (a) roll force and (b) torque per unit width during the rolling analysis step for implicit simulations with $N_{e}=5$, $N_{e}=30$ and $N_{e}=40$ CPE4R elements. Relatively steady average force and torque values are observed after $0.001\,$\mathrm{s}$$ for each simulation. All simulations have radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
Figure 4 plots (a) roll force and (b) torque per unit width as a function of simulation time during the rolling step for various mesh densities. We note that the $N_{e}=5$ simulation gives average roll force and torque values similar to the values outputted from the simulations with $N_{e}=30$ and $N_{e}=40$. Some authors use roll force and torque in their mesh convergence studies by identifying the mesh size at which no further significant changes in force or torque occur [62, 18]. By extension, we could take the $N_{e}=5$ results as sufficient here. However, as we will see in Figure 5, the simulation with $N_{e}=5$ performs poorly in predicting the von Mises stress, shear stress and plastic equivalent strain (PEEQ) values through the thickness of the sheet, thus proving that roll force and torque are poor measures of accuracy when through-thickness predictions are desired.
[…]
For von Mises stress, the largest and smallest numerical errors occur near $x/L=1.09$ (see the dashed lines in Figure 5(a)) and $x/L=0.6913$ (see the dotted lines in Figure 5(a)) respectively. This makes sense because the von Mises stress satisfies the yield condition uniformly in the interior of the contact region, except very close to the start and end where the material may be behaving elastically. For the shear stress, the largest errors occur at $x/L=0.025$ (see the solid curves in Figure 5(b)). The maximum relative difference in shear between the simulations with $N_{e}=30$ and $N_{e}=40$ is 4.9% at this position. Finally, the deviations in PEEQ between different mesh densities are largest near $x/L=1.09$ (see Figure 5(c)), but the PEEQ errors are much smaller than for von Mises stress and shear stress (see Table 4).
We note that the simulation with $N_{e}=5$ does not capture trends such as non-monotonic von Mises stress at $x/L=1.09$ (Figure 5(a)) or the curvature of the shear stress (Figure 5(b)). Yarita et al. [28] and Cawthorn et al. [39] used just $N_{e}=4$ and $N_{e}=5$ reduced-integration elements respectively through the half-thickness of the sheet, which our results suggest is far too coarse a mesh to accurately capture the through-thickness stress and strain distributions in the sheet. We use $N_{e}=30$ CPE4R elements through the half-thickness of the sheet in the remainder of this work since the maximum difference from the simulation with $N_{e}=40$ is 4.9%, and the computation time is 4.83 times faster compared to the $N_{e}=40$ simulation.
Figure 5: Interpolated (a) von Mises stress, (b) shear stress and (c) PEEQ (plastic equivalent strain) from simulations with varying mesh densities. The solid, dotted and dashed lines depict results in the sheet at $x/L=0.025$, $x/L=0.6913$ and $x/L=1.09$, respectively (see Figure 1). The horizontal centre (symmetry axis in Figure 1) of the sheet corresponds to $z/h_{0}=0$. All simulations employ a radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
Table 4: CPU time and maximum percentage error in von Mises stress, PEEQ (plastic equivalent strain) and shear stress for simulations with $N_{e}=$ 5–30 benchmarked against interpolated results from the simulation with $N_{e}=40$. Errors are calculated relative to the maximum absolute roll-gap value from the simulation with $N_{e}=40$. All simulations in this table are implicit simulations with CPE4R elements.
Elements through half-thickness CPU time (hours) Maximum von Mises % error Maximum shear % error Maximum PEEQ % error
5 0.25 20.55 45.48 7.24
10 1.16 12.97 24.3 4.33
15 3.2 9.74 14.51 2.93
20 7.09 5.71 9.24 2.02
30 15.89 3.68 4.9 1.09
40 76.75 - - -
4.4 Steady-state determination
Figure 6 plots interpolated von Mises stress, shear stress and PEEQ results from different time points throughout the $N_{e}=30$ simulation. Although Figure 4 suggests that the roll force and torque results reach a relatively steady value after time $t=0.001\,$\mathrm{s}$$, the interpolated results in Figure 6 suggest that the through-thickness stress and strain components are still significantly changing in time. This highlights the inadequacy of roll force and torque in terms of predicting the steady state of the rolling process. Despite this inadequate steady-state determination, roll force and torque are used by various authors [see, for example, 16, 43] for determining if a steady state has been reached by analysing the time at which these quantities become relatively constant. Although the roll force and torque are necessarily constant during the steady state of the rolling process, constant roll force and torque are not sufficient for identifying steady-state behaviour on their own, as shown here.
Figure 6 suggests the system is in a steady state by $t=0.0802\,$\mathrm{s}$$, since the through-thickness predictions do not significantly change from this time onwards. The data used for the remainder of this paper is therefore extracted at $t=0.0802\,$\mathrm{s}$$ to both ensure steady results and to facilitate the velocity calculation in the postprocessing (see Section 2 of the supplementary material).
Figure 6: Interpolated (a) von Mises stress, (b) shear stress and (c) PEEQ (plastic equivalent strain) from different time points of the implicit simulation with $N_{e}=30$ CPE4R elements. The solid, dotted and dashed lines depict results in the sheet at $x/L=0.025$, $x/L=0.6913$ and $x/L=1.09$, respectively (see Figure 1). The horizontal centre (symmetry axis in Figure 1) of the sheet corresponds to $z/h_{0}=0$. The simulation employs a radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
5 Comparison to analytical models
In this section, we compare FE simulations to the rigid-perfectly-plastic analytical model of Erfanian et al. [13]. We also compute the residual stress in the sheet after rolling and investigate its connection to the oscillatory pattern in stress and strain quantities in the roll gap. This connection is likely to have a strong relationship with curvature in asymmetric rolling, a phenomenon that is often incorrectly predicted [63].
5.1 Comparing FE outputs to the analytical results
Figure 7 shows shear stress contour plots ($\mathrm{M}\mathrm{P}\mathrm{a}$) comparing the FE outputs with the analytical results of Erfanian et al. [13]. Three roll-gap aspect ratios are investigated: $\varepsilon=0.175$, $\varepsilon=0.125$ and $\varepsilon=0.075$. As Erfanian et al. [13] do not consider deformations outside the roll gap, our attention here is restricted to values of $x/L$ between 0 and 1 (see Figure 1 for domain definition). For each value of $\varepsilon$, an oscillatory pattern is predicted by both the FE simulations and the analytical model. The number and position of the predicted shear stress lobes are in agreement for both methods. Similar to Minton [43], the number of lobes increases with decreasing $\varepsilon$. The magnitude of shear stress at various through-thickness positions is also noted to be, in general, similar. There are however two notable differences between the two approaches. Firstly, the lobes in the FE results have a more rounded profile compared to the analytical outputs, particularly for larger values of $\varepsilon$. As $\varepsilon$ decreases the lobe profiles become more similar. This is unsurprising as the analytical model of Erfanian et al. [13] is only valid in the limit $\varepsilon\rightarrow 0$.
$R=130.6\,$\mathrm{m}\mathrm{m}$$,
$\varepsilon=0.175$
$R=257.45\,$\mathrm{m}\mathrm{m}$$,
$\varepsilon=0.125$
$R=710\,$\mathrm{m}\mathrm{m}$$,
$\varepsilon=0.075$
Figure 7: Contour plots of shear stress results from simulations with $N_{e}=30$ elements. The left panel (a-c) shows FE simulations and the right panel (d-f) shows results from the analytical model of Erfanian et al. [13], for three different roll-gap aspect ratios: $\varepsilon=0.175$ (top row, $R=130.6\,$\mathrm{m}\mathrm{m}$$), $\varepsilon=0.125$ (middle row, $R=257.45\,$\mathrm{m}\mathrm{m}$$) and $\varepsilon=0.075$ (bottom row, $R=710\,$\mathrm{m}\mathrm{m}$$). Data is presented for the roll gap only, between $x/L=0$ and $x/L=1$, in all cases since Erfanian et al. [13] do not consider deformations outside the roll gap.
The second difference between the two approaches is the predicted shear profile between the neutral point and the end of the roll gap (i.e., $x_{N}<x/L<1$). The neutral point $x_{N}$ is defined as the position at which friction changes direction on the surface. In Figure 7 the neutral point is observed at $x/L\approx 0.6913$ in all contour plots. In the FE results, shear stress on the surface varies in a smooth fashion while the analytical model demonstrates a distinct discontinuity in shear stress. The discontinuous change in shear stress seen in the analytical model is due to the slipping Coulomb friction law that is employed. In reality, material sticks to the roller around the neutral point and shear stress changes sign smoothly [23, 14]. This sticking region around the neutral point is predicted in our FE results but not in the rigid-plastic analytical model.
Coulomb friction is employed in our FE model, but frictional constraints are enforced with the stiffness (penalty) method. The penalty method permits some relative surface motion (an “elastic slip”) when sticking should occur (i.e., when the shear stress is below the critical magnitude for sliding). While the surfaces are “sticking” the magnitude of sliding is limited to this elastic slip. Elastic slip is controlled computationally by a slip tolerance factor $\gamma$. In Figure 8 we investigate the effect that varying $\gamma$ has on the surface shear stress results, and observe that the gradient in surface shear steepens with decreasing $\gamma$, but to a finite limiting value that is far from the infinite slope predicted analytically.
Figure 8: Shear stress along the top surface of the deformed sheet from the analytical model by Erfanian et al. [13] and from simulations with different Young’s modulus ($E$) and elastic slip tolerance ($\gamma$) values. The default values are $E=206.3\,$\mathrm{G}\mathrm{P}\mathrm{a}$$ and $\gamma=0.005$. All simulations have radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
Figure 8 also shows that increasing the Young’s modulus by a factor of six has a significant effect on the numerically-predicted shear stress on the sheet’s surface: the gradient near the neutral point steepens and the position of the neutral point is closer to the position predicted analytically. This is expected as these material parameters more closely resemble rigid-plastic behaviour. This suggests that the friction rule alone is not responsible for the discrepancies between the FE and analytical results in Figure 7, and that the rigid-plastic (i.e., no elastic effects) assumption in the analytical model is also an important contributing factor.
Figure 9 shows von Mises stress and PEEQ rate contour plots, along with the sheet’s surface and central velocity magnitudes, from the two simulations with Young’s moduli $E=206.3\,$\mathrm{G}\mathrm{P}\mathrm{a}$$ and $E=1237.8\,$\mathrm{G}\mathrm{P}\mathrm{a}$$ ($6E$). The PEEQ rate results were calculated in a similar manner to the velocities (see Section 2 of the supplementary information provided). It is interesting to note the clearly non-vertical boundaries between the sub-yield regions and the yielding regions around the roll gap entrance and exit in Figures 9(a) and 9(d). Many analytical models make the assumption of vertical boundaries between these regions. Another interesting point to note is that the end of the plastic deformation zone is closer to $x/L=1$ in Figure 9(d) than in Figure 9(a), indicating less elastic springback for a higher Young’s modulus. This is unsurprising since elastic strains decrease as $E\rightarrow\infty$.
There are differences between Figures 9(b) and 9(e), the two PEEQ rate contour plots, but the general trends are the same. A clear oscillatory pattern exists in the PEEQ rate plots, and regions of zero-valued PEEQ rate periodically appear at the sheet’s surface and centre. Regions of zero PEEQ rate indicate that no plastic flow is occurring in these regions. However, the von Mises results do not indicate that the stress is significantly below the yielding value in the neighbourhood of any of these no-plastic-flow zones. Therefore we refer to these regions as no-plastic-flow regions as opposed to “sub-yield” regions. The surface and central velocity plots in Figures 9(c) and 9(f) also show clear oscillatory behaviours. Each region of constant velocity at the sheet’s surface or centre coincides with a no-plastic-flow region at the corresponding sheet location. This is plausible since constant velocity implies a rigid-body motion. We also see a sticking region in both surface velocity outputs around the neutral point, where the speed of the surface of the sheet is equal to the roller’s tangential speed. This allows the shear stress to smoothly change from positive to negative on the surface as the sheet sticks to the rollers for a small portion of the roll gap. A smaller sticking region is observed for the $6E$ simulation. We hypothesise that this sticking region would continue to shrink, and hence tend toward the analytical output, in the limit as $E\rightarrow\infty$, but given convergence issues for higher $E$ values than are given here, this remains a hypothesis.
We have shown that by decreasing the value of $\gamma$ and increasing the value of $E$, the FE results become more similar to the analytical results of Erfanian et al. [13]. Therefore, we hypothesise that by taking the limit as $\gamma\rightarrow 0$ and as $E\rightarrow\infty$, it would be possible to achieve much better agreement between the FE results and the rigid-plastic shear stress results given by Erfanian et al. [13] around the neutral point. However, decreasing the slip tolerance $\gamma$ and increasing the Young’s modulus $E$ are computationally intensive tasks that suffer from convergence issues when more extreme values are attempted. The limitations of the analytical model of Erfanian et al. [13] in terms of accurately predicting the shear stress distribution are therefore clear: (1) The shear lobe shape deviates from FE predictions as the roll-gap aspect ratio $\varepsilon$ increases from zero and (2) the material and friction models are inadequate for capturing sticking behaviours and the smoothly-varying shear stress around the neutral point.
Figure 9: Comparison of FE simulations for two Young’s moduli. The left panel (a–c) shows results from the FE simulation with the original Young’s modulus ($E=206.3\,$\mathrm{G}\mathrm{P}\mathrm{a}$$). The right panel (d–f) shows results for a Young’s modulus six times greater than that of (a–c). Panels (a) and (d) show the von Mises contour plots for each simulation, while panels (b) and (e) show the PEEQ rate contour plots. Panels (c) and (f) show the sheet’s surface and central velocity magnitudes (solid and dashed lines respectively) for each simulation, and the roller’s tangential velocity (horizontal black dotted line). The PEEQ rate and velocity values are obtained through postprocessing of the PEEQ values and displacements respectively (see Section 2 of the supplementary material). Both simulations employ $N_{e}=30$ CPE4R elements, a radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
5.2 Residual stress
Here we display the residual stress in the sheet after rolling and investigate its connection to the through-thickness variation of the shear stress in the roll gap. To highlight this connection, we consider a symmetric simulation with a roller radius $R=32\,$\mathrm{m}\mathrm{m}$$, which leads to a roll-gap aspect ratio of $\varepsilon=0.354$, and all other parameters are the same as described in Section 3. We measure the von Mises stress and the quantities $p\pm 2\theta Y/\sqrt{3}$, where $p=-0.5\left(\sigma_{xx}+\sigma_{zz}\right)$ is the hydrostatic pressure and $\theta$ is the anticlockwise rotation angle of the stress element from the positive $x$-axis for which the stress is in a state of pure shear. The von Mises stress in the sheet after it passes through the roll gap can be used to give an indication of residual stress, while the quantities $p\pm 2\theta Y/\sqrt{3}$ are of interest because these quantities are constant along two orthogonal sets of characteristic lines for plastic deformation, and the direction of these characteristic lines coincide with the lines of maximum shear stress (slip lines) when plastic deformation is occurring (see Section 5 of the supplementary material for more details).
Figure 10 shows a contour plot of von Mises stress and of $p\pm 2\theta Y/\sqrt{3}$ ($\mathrm{M}\mathrm{P}\mathrm{a}$) from a symmetric simulation with $N_{e}=30$. The right-end $x/L$ value in the plot is large and the von Mises stress towards the right end of the contour plot can therefore give an indication of residual stress since it is significantly far from the external loading (i.e., the rollers). The distinct centre and outer von Mises zones in the sheet after the roll gap in Figure 10(a) correspond to zones of lengthwise compression and tension respectively. This is consistent with longitudinal stress results reported by Tadić et al. [45]. The length–height ratio of these contour plots is set to the true ratio used in the simulation. This allows for interpretation of the direction of the contour lines in Figure 10(b). For large portions of the roll gap, these contour lines meet each other at $90^{\circ}$, indicating that plastic deformation is occurring. The sub-yield are easily identified by comparing against the von Mises contour plot in Figure 10(a). Recall the initial yield stress is $Y=477.2\,$\mathrm{M}\mathrm{P}\mathrm{a}$$, and thus the regions of the sheet with von Mises values significantly less than this value are sub-yield regions. In these sub-yield regions the $p\pm 2\theta Y/\sqrt{3}$ contour lines are not orthogonal, and thus cannot be called slip lines in these regions. The contour lines are only slip lines in regions of plastic deformation. Another interesting observation from the von Mises contour plot in Figure 10(a) is that there is a significant amount of residual stress under the top surface (and above the bottom surface) of the sheet for $x/L>1$. The $p\pm 2\theta Y/\sqrt{3}$ contour lines oscillate through the roll gap in a similar manner to the shear lobes discussed earlier, which is unsurprising given they line up with the directions of maximum shear stress during plastic flow. After oscillating through the roll gap, the $p\pm 2\theta Y/\sqrt{3}$ contour lines also accumulate within a thin layer that roughly lines up with these regions of residual von Mises stress. This indicates that a strong relationship exists between the oscillations in shear stress in the roll gap and residual stress outside the roll gap.
Figure 10: Contour plots of (a) von Mises and (b) $p\pm 2\theta Y/\sqrt{3}$, all given in $\mathrm{M}\mathrm{P}\mathrm{a}$, from a symmetric implicit simulation with $N_{e}=30$ CPE4R elements. The radius is $R=32\,$\mathrm{m}\mathrm{m}$$ which gives a roll-gap aspect ratio of $\varepsilon=0.354$. The tangential velocity of the rollers is $R\Omega=1287.25\,$\mathrm{m}\mathrm{m}\,\mathrm{s}^{-1}$$.
5.3 Asymmetric rolling
We propose that the connection between shear stress in the roll gap and residual stress significantly far away from the roll gap can be used for curvature prediction in asymmetric rolling. For this investigation, we consider an asymmetric simulation with a roller radius $R=32\,$\mathrm{m}\mathrm{m}$$, where the top roller rotates 2.5% faster than the bottom roller. The bottom roller rotates with the same tangential velocity as before ($R\Omega=1287.25\,$\mathrm{m}\mathrm{m}\,\mathrm{s}^{-1}$$), and all other parameters are the same as described in Section 3.
The asymmetric results shown in Figure 11 are obtained via a simulation with $402N_{e}=$ 40. We use this mesh density here as the equivalent of the mesh size used in previous simulations ($602N_{e}=$ 60) was unachievable due to the whole sheet having to be discretised in the asymmetric simulation as opposed to just the top half. Large amounts of residual stress are observed in the asymmetric von Mises results in Figure 11(a), similar to what was noted in the symmetric case (Figure 10(a)). Once again the contour lines of $p\pm 2\theta Y/\sqrt{3}$ accumulate within a thin layer that lines up with these regions of residual von Mises stress. The magnitude of the residual stress in the sheet for $x/L>1$ is greater towards the bottom of the sheet, and the sheet’s curvature is towards the slower roller for this specific set of parameters. We do not suggest that curvature is always towards the slower roller, because as pointed out by Minton and Brambley [63], many authors have given contradicting prediction rules for curvature during asymmetric rolling. Rather we hypothesise that the oscillatory pattern in stress and strain quantities presented here, which is largely absent in the literature, has a strong influence on the accurate prediction of the residual stress for $x/L>1$, which may have a significant impact on predicting curvature in asymmetric rolling. This is an important aspect of modelling rolling since unintentional asymmetries cause turn-up and turn-down in sheet rolling which can halt the process or even damage the rolling table and mills and since asymmetries are intentionally exploited in asymmetrical rolling to improve process efficiency and produce curved products [63].
Figure 11: Contour plots of (a) von Mises and (b) $p\pm 2\theta Y/\sqrt{3}$, given in $\mathrm{M}\mathrm{P}\mathrm{a}$, from an implicit simulation with $402N_{e}=$ 40 CPE4R elements and asymmetric roller speeds. The radius is $R=32\,$\mathrm{m}\mathrm{m}$$ which gives a roll-gap aspect ratio of $\varepsilon=0.354$. The top roller rotates 2.5% faster than the bottom roller. The tangential velocity of the bottom roller is $R\Omega=1287.25\,$\mathrm{m}\mathrm{m}\,\mathrm{s}^{-1}$$.
6 Conclusion
[…]
We show that our through-thickness predictions compare well with a rigid-perfectly-plastic analytical model [13]. When optimized correctly, analytical mathematical models can be used in online control due to their rapid computation time and ability to provide accurate results compared to FE simulations. Excellent agreement in shear stress was found when comparing the analytical and FE models. Both models captured the oscillatory nature of shear during rolling. The relatively small discrepancies between the shear results of the two models can be attributed to the size of the roll-gap aspect ratio, $\varepsilon$ and the differences in the material and friction models. In the limits as $\varepsilon\rightarrow 0$, $\gamma\rightarrow 0$ and $E\rightarrow\infty$, we expect that the agreement between the FE and analytical results would be even further improved.
Without the ability to compare to an analytical model, previous studies have often relied on the convergence of roll force and roll torque to assess the quality of numerical rolling simulations. Here, we show that roll force and roll torque are poor indicators of simulation quality. This is because these averaged quantities often converge well (see Figure 4(b)) even for under-resolved simulations that fail to accurately predict other quantities of interest (such as the distribution of shear stress at the roll exit in Figure 5(b)). Comparison to a steady state analytical model does however require that time-domain FE simulations converge to a steady state. Here, instead of starting our simulations with the sheet in front of the rollers and pushing the sheet into the roll gap, as is common practice, we start with the sheet in the roll gap but with the rollers removed and then squeezed the rollers into the sheet in a preliminary bite step before rolling (as described in subsection 3.3). This was found to improve convergence to a steady state of rolling.
Unexpectedly, our results show the existence of a region near the neutral point where the sheet and rollers are sticking, rather than slipping. This is not commonly described in previous studies, even in studies of extreme cases such as foil rolling where many different regions within the roll gap are detailed and analysed (e.g., Fleck et al. [64], Figure 3c, zones E and F). The assumption that only slipping occurs at the neutral point is a common one, and indeed it is the cause of the discrepancy between the present FE model and the analytical model. While care has been taken here to ensure that the sticking region near the neutral point is not a numerical artefact, it is a consequence of Coulomb friction, which may be an overly-simplified model of the friction that practically occurs in metal rolling. Comparison with the analytical model shows that the effects of this sticking region on the stress oscillation pattern through-thickness are important beyond the neutral point and affects the shear stresses at the roll-gap exit.
We also investigated the effect of the oscillatory pattern in shear stress on residual stress. We found that a strong relationship exists between regions of large residual stress on the exit side of the roll gap and the $p\pm 2\theta Y/\sqrt{3}$ contour lines, whose direction coincides with the directions of maximum shear during plastic flow. Since the oscillatory pattern is largely absent from the existing analytical and numerical literature, this pattern provides a promising route to curvature predictions in asymmetric rolling.
Future work could include employing second-order elements to discretise the metal sheet or incorporating the elastic deformation of the work roller. Elastic roll deformation would be particularly important for foil rolling, where the roll-gap aspect ratio is very large and the through-thickness oscillatory pattern will repeat many times; the consequences of these oscillations in foil rolling are currently unknown. Extending the preliminary study in subsection 5.3 of curvature induced by asymmetric rolling might allow for the prediction and control of curvature; the lack of through-thickness resolution may be what has been preventing previous FE studies from correctly predicting curvature from asymmetric rolling [63]. While in this study the FE software ABAQUS was used, we believe the principles described here are applicable to all FE studies, and this could be verified and best-practice refined by performing similar studies using other commonly-used software. Finally, given the detailed preparation of this particular FE model, we are confident that the final FE model described in this paper can help to guide the formulation of simpler, faster mathematical models with wave-like solutions.
Acknowledgements
This publication has emanated from research conducted with the financial support of Science Foundation Ireland under grant number 18/CRT/6049. For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
AOC is supported by the European Union, Science Foundation of Ireland and Lero, the Science Foundation Ireland Research Centre for Software, grants #101028291 #13/RC/2094 and #SOWP2-TP0023, respectively. ME gratefully acknowledges the funding of a University of Warwick Chancellor’s Scholarship. EJB is grateful for the UKRI Future Leaders’ Fellowship funding (MR/V02261X/1) supporting this work. DOK is grateful for funding from Science Foundation Ireland (21/FFP-P/10160).
Supplementary Material
Appendix S1 ABAQUS formulation
We present here some relevant background information on the FE modelling technique for rolling. The metal rolling process is governed by differential equations describing yield conditions, constitutive relationships and conservation of mass and momentum. In FE analysis, the differential equations are approximately solved between a discrete set of nodes, using shape functions that assume a particular interpolation field between the nodes. This approach discretises spatial derivatives, so that the partial differential equations are transformed into a system of ordinary differential equations. The simplified form of the equations of motion are written as
$\mathbf{M}\ddot{\mathbf{u}}=\mathbf{f}^{\text{ext}}-\mathbf{f}^{\text{int}},$ (1)
where $\mathbf{M}$ is the diagonal mass matrix, $\mathbf{u}$ is the nodal displacement vector, and $\mathbf{f}^{\text{ext}}$ and $\mathbf{f}^{\text{int}}$ are the external and internal nodal force vectors respectively [57]. In metal rolling, inertia effects are negligible and are therefore often ignored [65, 14, 43]. The rolling process can be idealised by assuming the stresses in the sheet are in quasi-static equilibrium. The implicit and explicit solvers offered by ABAQUS are both useful for obtaining solutions to FE problems concerning quasi-static processes. We now give a brief overview of the formulations of ABAQUS/Standard and ABAQUS/Explicit.
S1.1 ABAQUS/Standard formulation
ABAQUS/Standard is an implicit solver where the model is updated from time $t$ to time $t+\Delta t$ using information from times $t$ and $t+\Delta t$. In this work, only static analyses are considered in ABAQUS/Standard. Therefore, since inertia is ignored, the equations of motion are now
$\mathbf{F}(\mathbf{u})=\mathbf{f}^{\text{ext}}-\mathbf{f}^{\text{int}}=\mathbf{0}.$ (2)
This large system of equations is solved iteratively in each time increment to obtain the updated nodal displacements, using the Newton–Raphson technique. The displacements are updated via
${\Delta\mathbf{u}_{\left(n+1\right)}}^{\left(t+\Delta t\right)}=\left[{\mathbf{K}_{\left(n\right)}}^{\left(t+\Delta t\right)}\right]^{-1}\cdot\left({\mathbf{F}_{\left(n\right)}}^{\left(t+\Delta t\right)}\right),$ (3)
${\mathbf{u}_{\left(n+1\right)}}^{\left(t+\Delta t\right)}={{\mathbf{u}}_{(n)}}^{\left(t+\Delta t\right)}+\Delta{\mathbf{u}_{\left(n+1\right)}}^{\left(t+\Delta t\right)},$ (4)
where the superscript $\left(t+\Delta t\right)$ is the time at which the equations are evaluated, the subscripts $\left(n\right)$ and $\left(n+1\right)$ indicate the current and next iterations respectively and $\mathbf{K}_{\left(n\right)}=\left(\partial\mathbf{F}/\partial\mathbf{u}\right)_{\left(n\right)}$ is the stiffness matrix at iteration $n$. The inversion of the stiffness matrix $\mathbf{K}_{\left(n\right)}$ is computationally expensive. The convergence of the Newton–Raphson method is measured by ensuring that all entries $\mathbf{F}(\mathbf{u})$ and $\Delta\mathbf{u}$ are sufficiently small [66]. At the end of time increment $t$, the next time increment $t+\Delta t$ is attempted, and load or displacement is incremented corresponding to $\Delta t$. If convergence is not achieved by ABAQUS in a certain time increment, $\Delta t$ is reduced and the Newton-Raphson method is attempted again. The reader should note that no velocities or accelerations are given by ABAQUS/Standard for static analyses. These can be manually extracted via postprocessing of the displacements (see Section S2).
S1.2 ABAQUS/Explicit formulation
In ABAQUS/Explicit, the equations are solved for future time steps using known information, making it an explicit scheme. This dynamic solution scheme necessitates the full version of equation (1) to be solved. The diagonal mass matrix is inverted to find the vector of nodal accelerations,
$\ddot{\mathbf{u}}^{(i)}=\mathbf{M}^{-1}\left(\mathbf{f}^{\text{ext}}-\mathbf{f}^{\text{int}}\right),$ (5)
at the beginning of time increment number $i$. The inversion of the mass matrix is numerically much simpler than inverting the stiffness matrix [67]. The acceleration vector $\ddot{\mathbf{u}}^{(i)}$ is assumed to remain constant for a very small time increment and the explicit central-difference integration rule [18] is used to update the mid-increment velocities and subsequently the end-of-increment displacements,
$\dot{\mathbf{u}}^{\left(i+\frac{1}{2}\right)}=\dot{\mathbf{u}}^{\left(i-\frac{1}{2}\right)}+\frac{\Delta t^{\left(i+1\right)}+\Delta t^{\left(i\right)}}{2}\ddot{\mathbf{u}}^{(i)},$ (6)
${\mathbf{u}}^{\left(i+1\right)}={\mathbf{u}}^{\left(i\right)}+\Delta t^{\left(i+1\right)}\dot{\mathbf{u}}^{\left(i+\frac{1}{2}\right)}.$ (7)
The explicit dynamics procedure is ideally suited for analysing high-speed dynamic events, but many of the advantages of the explicit procedure also apply to the analysis of slower (quasi-static) processes. Quasi-static analyses typically have longer timescales which increase the numerical computation time required by FE methods. The stable time increment for ABAQUS/Explicit simulations is the largest possible time step the solver can take without results becoming unstable. When the solution becomes unstable, the time-history response of solution variables such as displacements will usually oscillate with increasing amplitudes [15]. This stability limit is also known as the Courant–Friedrich–Lewy (CFL) limit [19], and is estimated as
$\Delta t_{\text{stable}}\approx\frac{L^{e}\sqrt{\rho}}{\sqrt{E}},$ (8)
where $L^{e}$ is the smallest element dimension in the FE model, $\rho$ is the material density and $E$ is the Young’s modulus of the material. However, ABAQUS/Explicit enables the use of mass scaling. Mass scaling uniformly scales the density of the material within the simulation to facilitate longer time steps and reduce computational time [43]. The limited time increment means that explicit methods are conditionally stable, thus the mass scaling factor (MSF) must be chosen carefully to reduce simulation times without introducing inertial effects. The introduction of inertial effects can lead to erroneous and unstable outputs [68, 69, 57]. The general rule of thumb to assess inertial effects is to ensure that the ratio between the kinetic energy and the internal energy is small (5-10%) [67, 15].
Appendix S2 FE velocity calculation
ABAQUS/Standard does not output velocity components or PEEQ rate values for static analyses. Therefore, we manually calculate the velocity and PEEQ rate values used to produce Figure 9. The velocity and PEEQ rate approximations are conducted via a Lagrangian time derivative of the displacements and the PEEQ values respectively. Since the procedure is identical for calculating both velocity and PEEQ rate values, we give details of the velocity calculation only.
All stress and strain quantities are outputted to an output database (.ODB) file for all ABAQUS simulations. These quantities are available at different time points throughout the simulation, and it is possible for the user to vary the number and frequency of the time points throughout each analysis step. To approximate the velocity components $\left(v_{x},v_{z}\right)$ of some arbitrary node in the sheet, at time frame $f$, we use a fourth-order-accurate central differencing operator on the displacements
${v}^{(f)}_{x}=\frac{\left(\frac{1}{12}\right){u}^{(f-2)}_{x}-\left(\frac{2}{3}\right){u}^{(f-1)}_{x}+\left(\frac{2}{3}\right){u}^{(f+1)}_{x}-\left(\frac{1}{12}\right){u}^{(f+2)}_{x}}{\Delta t},$ (9a)
${v}^{(f)}_{z}=\frac{\left(\frac{1}{12}\right){u}^{(f-2)}_{z}-\left(\frac{2}{3}\right){u}^{(f-1)}_{z}+\left(\frac{2}{3}\right){u}^{(f+1)}_{z}-\left(\frac{1}{12}\right){u}^{(f+2)}_{z}}{\Delta t},$ (9b)
where $\Delta t$ is constant between all time frames considered. In order to allow velocities and all other stress and strain components to be evaluated at the same time frame, all data in Section 5 is evaluated at a time point that is two time frames before $t=0.1\,$\mathrm{s}$$, which is the end of the rolling step. The authors ensured that $\Delta t$ is sufficiently small for accurate velocity approximations.
Appendix S3 Incompatible CPE4I elements
Figure S1 compares interpolated shear stress from two simulations: one with CPE4I elements and the other with CPE4R elements. Both simulations employ $N_{e}=15$ elements through the half-thickness, since simulations with $N_{e}\geq 20$ CPE4I elements required extremely small time increments which gave rise to convergence issues. The maximum relative error in Figure S1 is approximately 9% when compared to the maximum absolute shear stress in the roll gap. We expect that this error would reduce significantly as the mesh density increases. Given the expected similar shear stresses with increasing mesh density and that the simulation with CPE4I elements requires 17.7 CPU hours compared to 3.2 CPU hours for the simulation with CPE4R elements, CPE4I elements were not used for the main results in this work.
Figure S1: Interpolated results for shear stress ($\mathrm{M}\mathrm{P}\mathrm{a}$) from simulations with CPE4I elements and CPE4R elements. Both simulations employ $N_{e}=15$ elements, a radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$. The results are almost identical and the CPU time is 17.7 hours for the CPE4I simulation compared to 3.2 hours for the CPE4R simulation. The solid, dotted and dashed lines depict the interpolated results in the sheet at $x/L=0.025$, $x/L=0.6913$ and $x/L=1.09$ respectively (see Figure 1). The horizontal centre (symmetry axis in Figure 1) of the sheet corresponds to $z/h_{0}=0$.
Appendix S4 Relative error: choice of $x/L$ locations
We detail here how the three $x/L$ positions used in the interpolated results throughout this paper were chosen. These $x/L$ positions represent the positions of maximum relative errors. To explain how these errors are calculated, let us consider the solid-line shear results in Figure 5b. The interpolated shear result from the simulation with $N_{e}=5$ is interpolated onto a curve with 40 points. The absolute difference between this new curve and the interpolated shear result from the simulation with $N_{e}=40$ is calculated at each $z/h_{0}$ point. Finally, the absolute difference at each point is divided by the maximum absolute shear stress in the entire roll gap, as predicted by the simulation with $N_{e}=40$. At each $x/L$ location, the maximum relative error value out of the 40 values is stored in an array. The same procedure is carried out for $N_{e}=10$, $N_{e}=15$, $N_{e}=20$ and $N_{e}=30$, and all of these maximum errors are plotted in Figure S2.
Figure S2: Relative error in (a) von Mises stress, (b) shear stress and (c) PEEQ (plastic equivalent strain) from simulations with varying mesh densities. All simulations are compared to a simulation with $N_{e}=40$ elements through the half-thickness. The error value at each $x/L$ location represents the maximum error through the thickness of the sheet at that $x/L$ position. All simulations employ a radius $R=257.45\,$\mathrm{m}\mathrm{m}$$ and an aspect ratio of $\varepsilon=0.125$.
We considered 100 $x/L$ positions from $x/L=0.025$ up to $x/L=1.25$. Note that $x/L<0.025$ was not considered. The contact point occurs in the neighbourhood of $x/L=0$ and corresponds to a change from zero to large normal stress, so a small change in the predicted position of the contact point can lead to misleadingly large errors recorded at $x/L=0$.
Figure S2 shows that $x/L=0.025$ is a position where high relative errors occur, and hence this is one of the $x/L$ positions used for interpolations in this paper. Figure 2(a) and Figure 2(b) also show a spike in relative errors near $x/L=1.09$. This is another $x/L$ position used to calculate the interpolated results. It is not clear from Figure 2(b) and Figure 2(c) which third $x/L$ position to choose. However, Figure 2(a) shows a small peak in relative error around the neutral point $x/L=0.6913$ (as predicted by the implicit simulation with $N_{e}=40$). Given this peak and the interesting dynamics around the neutral point in general, $x/L=0.6913$ is the third position used in the interpolated results throughout this paper.
Appendix S5 Slip line theory
This theory is only applicable to rigid-perfectly-plastic materials. In two-dimensional plane-strain analyses, the stress tensor takes the form
$\boldsymbol{\sigma}$ $=\boldsymbol{S}-p\boldsymbol{I}=\begin{pmatrix}S_{xx}&S_{xz}\\ S_{xz}&S_{zz}\end{pmatrix}+\begin{pmatrix}-p&0\\ 0&-p\end{pmatrix}$ (10)
$\phantom{=}\quad\quad\quad\,=\begin{pmatrix}S_{xx}-p&S_{xz}\\ S_{xz}&S_{zz}-p\end{pmatrix},$
where $\boldsymbol{S}$ is the deviatoric part of the stress tensor and $-p=\left(\sigma_{xx}+\sigma_{zz}\right)/2$ is the two-dimensional hydrostatic pressure [48]. Under rotation of the coordinate system by $\theta$ degrees from the positive $x$-axis in the anticlockwise direction, the stress state $\boldsymbol{\sigma}$ may be rewritten as
$\boldsymbol{\sigma}^{\prime}$ $=\boldsymbol{R}\boldsymbol{\sigma}\boldsymbol{R}^{T}$ $=\begin{pmatrix}\cos{\theta}&\sin{\theta}\\ -\sin{\theta}&\cos{\theta}\end{pmatrix}\begin{pmatrix}S_{xx}-p&S_{xz}\\ S_{xz}&S_{zz}-p\end{pmatrix}\begin{pmatrix}\cos{\theta}&-\sin{\theta}\\ \sin{\theta}&\cos{\theta}\end{pmatrix}$ (11)
$=\begin{pmatrix}S_{xx}\cos{2\theta}+S_{xz}\sin{2\theta}-p&S_{xz}\cos{2\theta}-S_{xx}\sin{2\theta}\\ S_{xz}\cos{2\theta}-S_{xx}\sin{2\theta}&-S_{xx}\cos{2\theta}-S_{xz}\sin{2\theta}-p\end{pmatrix},$
where both $\boldsymbol{\sigma}$ and $\boldsymbol{\sigma}^{\prime}$ represent the same stress state in different coordinate systems, and $\boldsymbol{R}$ is a rotation matrix [70]. It is possible to write the von Mises yield criterion in terms of the deviatoric stress components as $S_{ij}S_{ij}=2Y^{2}/3$, where $Y$ is the yield stress in tension or compression of the material. For some rotation angle $\theta$, the stress state $\boldsymbol{\sigma}^{\prime}$ is in a state of pure shear,
$\boldsymbol{\sigma}^{\prime}$ $=\begin{pmatrix}-p&\frac{Y}{\sqrt{3}}\\ \frac{Y}{\sqrt{3}}&-p\end{pmatrix},$ (12)
and the stress state $\boldsymbol{\sigma}$ in the original coordinate system can be written as
$\boldsymbol{\sigma}$ $=\begin{pmatrix}p-\frac{Y}{\sqrt{3}}\sin{2\theta}&\frac{Y}{\sqrt{3}}\cos{2\theta}\\ \frac{Y}{\sqrt{3}}\cos{2\theta}&p+\frac{Y}{\sqrt{3}}\sin{2\theta}\end{pmatrix}.$ (13)
We can use this definition for $\boldsymbol{\sigma}$ in the two-dimensional stress equilibrium equations to write
$0=\frac{\partial\sigma_{xx}}{\partial x}+\frac{\partial\sigma_{xz}}{\partial z}=-\frac{\partial p}{\partial x}-\frac{2Y}{\sqrt{3}}\left(\cos{2\theta}\frac{\partial\theta}{\partial x}+\sin{2\theta}\frac{\partial\theta}{\partial z}\right),$ (14)
$0=\frac{\partial\sigma_{xz}}{\partial x}+\frac{\partial\sigma_{zz}}{\partial z}=-\frac{\partial p}{\partial z}-\frac{2Y}{\sqrt{3}}\left(\sin{2\theta}\frac{\partial\theta}{\partial x}-\cos{2\theta}\frac{\partial\theta}{\partial z}\right).$ (15)
Equations 14 and 15 are hyperbolic PDEs and can be solved by the method of characteristics, (see Johnson et al. [47] for a detailed description). The characteristic directions of each of these equations are given by
$\dfrac{dz}{dx}=\tan{\theta},$ (16)
$\dfrac{dz}{dx}=\tan{\left(\theta+\frac{\pi}{2}\right)},$ (17)
respectively, and it is clear that the families of characteristics form an orthogonal network. The family of characteristics given by the $\theta$ parameter are called the $\alpha$-lines and those given by the $\theta+\pi/2$ parameter are called $\beta$-lines. Both sets of lines coincide with the lines of maximum shear stress. It can be shown that $p+2\theta Y/\sqrt{3}$ is constant along the $\alpha$-lines and $p-2\theta Y/\sqrt{3}$ is constant along the $\beta$-lines. However, these quantities are only constant for material that has yielded plastically [47]. Similarly, the lines will not be orthogonal in regions where the material is not deforming plastically.